### v2

# Betti numbers of the moduli space of Higgs bundles over a real curve

Abstract.
We produce a formula for the $\mathbb{Z}_{2}$-Betti numbers of the moduli space $M_{r}^{d}$ of stable real Higgs bundles over a real projective curve, with coprime rank $r$ and degree $d$. Our approach relies on the motivic formula for the moduli space due to Mellit [M], Fedorov-Soibelman-Soibelman [FSS], and Schiffmann [S], and the fact that the virtual $\mathbb{Z}_{2}$ Poincaré polynomial is a motivic measure over $\mathbb{R}$.
Key words and phrases:
Moduli spaces of Higgs bundles, symmetric products of a curve, real algebraic geometry, motivic ring, Poincaré polynomials, Plethystic exponential
1. Introduction
Let $\Sigma/k$ be a smooth, projective, geometrically connected curve of genus $g$ over a field $k$ admitting a $k$-rational point, and denote by $M_{r}^{d}$ the moduli space of stable Higgs bundles of rank $r$ and degree $d$ over $\Sigma$. When $gcd(d,r)=1$, which we assume henceforth, $M_{r}^{d}$ is a smooth, quasi-projective variety.
For $g,r,d\in\mathbb{Z}$ and $g,r\geq 0$, Schiffmann [S] defines a Laurent polynomial
$A_{g,r,d}(q,\alpha_{1},...,\alpha_{g})\in\mathbb{Z}[q,\alpha_{1}^{\pm 1},...,\alpha_{g}^{\pm 1}]$
which is symmetric under permutations of the $\alpha_{i}$ and under $\alpha_{i}\mapsto q\alpha_{i}^{-1}$. If $k=\mathbb{F}_{q}$ and $\{\alpha_{i},\alpha_{i}^{-1}q|i=1,..,g\}$ are set to the eigenvalues of the Frobenius action on $H^{1}_{et}(\Sigma,\mathbb{Q}_{l})$ for $(q,l)=1$, then $A_{g,r,d}$ counts indecomposable bundles of rank $r$ and degree $d$ over $\Sigma$, and $q^{1+(g-1)r^{2}}A_{g,r,d}$ counts points in $M_{r}^{d}(\mathbb{F}_{q})$. Using the Artin-Grothendieck Comparision Theorem, it follows that that when $k=\mathbb{C}$, the Poincaré polynomial for the compactly supported rational Betti numbers of $M_{r}^{d}(\mathbb{C})$ equals
$\sum_{k}(-1)^{k}t^{k}\dim_{\mathbb{Q}}H^{k}_{cpt}(M_{r}^{d}(\mathbb{C});\mathbb{Q})=t^{2(g-1)r^{2}+2}A_{g,r,d}(t^{2},t,t,t,...,t,t)$
Our main result is that a similar formula calculates $\mathbb{Z}_{2}$ Betti numbers of $M_{r}^{d}(\mathbb{R})$ when $k=\mathbb{R}$. For a variety $X/\mathbb{R}$, denote its compactly supported Poincaré polynomial by
$P_{t}(X):=\sum_{k}(-1)^{k}t^{k}\dim_{\mathbb{Z}_{2}}H^{k}_{cpt}(M_{r}^{d}(\mathbb{R});\mathbb{Z}_{2}).$
Theorem 1.1.
Suppose that $\Sigma$ is a smooth, projective, geometrically connected curve of genus $g$ defined over $k=\mathbb{R}$, and suppose that $\Sigma(\mathbb{R})$ is a union of $b+1$ circles, with $b\geq 0$. Then
$P_{t}(M_{r}^{d})=(-t)^{(g-1)r^{2}+1}A_{g,r,d}(-t,-t^{\frac{1}{2}},-t^{\frac{1}{2}},...,-t^{\frac{1}{2}},-1,-1,..,-1)$
where (gb)(g-b)-many entries equal $t^{\frac{1}{2}}$ and $b$-many equal 1-1.
For example, for $r=2$, we obtain the explicit formula
$P_{t}(M_{2}^{1})=2^{b}t^{4(g-1)+1}(1-t)^{g}f(t^{1/2},1)$
where
$\begin{split}f(t^{\frac{1}{2}},z)&=\frac{(z^{2}+t^{\frac{1}{2}})^{g-b}(z-t^{\frac{1}{2}})^{g-b}(z^{2}+1)^{b}(z-t)^{b}}{(z^{2}-1)(z+t)}+\frac{(z-t^{\frac{3}{2}})^{g-b}(1+t^{\frac{3}{2}})^{g-b}(z-t)^{b}(1+t^{2})^{b}}{(z+t)(1-t^{2})}\\ &-\frac{(z+t^{\frac{1}{2}})^{g-b}(1-t^{\frac{1}{2}})^{g-b}(z+1)^{b}(1-t)^{b}}{2(z-1)(1+t)}-\frac{(z-t^{\frac{1}{2}})^{g-b}(1+t^{\frac{1}{2}})^{g-b}(z-1)^{b}(1+t)^{b}}{2(z+1)(1-t)}.\end{split}$
A different formula for $P_{t}(M_{2}^{1})$ was obtained in [B]. Formulas for $P_{t}(M_{r}^{d})$ when $r\geq 3$ are new.
We use a construction of $A_{g,r,d}$ due to Mellit [M]. Let $\mathcal{P}$ denote the set of Young diagrams. Given $\mu\in\mathcal{P}$, define the rational function $\mathcal{H}_{\mu}$ in variables $q,z,\alpha_{1},...,\alpha_{g}$,
$\mathcal{H}_{\mu}:=\prod_{\square\in\mu}\frac{\prod_{i=1}^{g}(z^{a(\square)+1}-\alpha_{i}q^{l(\square)})(z^{a(\square)}-\alpha_{i}^{-1}q^{l(\square)+1})}{(z^{a(\square)+1}-q^{l(\square)})(z^{a(\square)}-q^{l(\square)+1})}$
where $a(\square)$ and $l(\square)$ denote the arm and leg lengths of boxes in the diagram. For $g\geq 1$, define rational functions $H_{g,r}$ by
(1.1) $\sum_{r=1}^{\infty}H_{g,r}T^{r}=(z-1)(1-q)\operatorname{Log}\sum_{\mu\in\mathcal{P}}T^{|\mu|}\mathcal{H}_{\mu},$
where $\operatorname{Log}$ denotes the plethystic logarithm. Mellit proved that for all $r\geq 1$, $H_{g,r}$ is a Laurent polynomial in $q$, $z$, and $\alpha_{1},\ldots,\alpha_{g}$, and for all $d$,
$A_{g,r,d}(q,\alpha_{1},\ldots,\alpha_{g})=H_{g,r}(q,1,\alpha_{1},\ldots,\alpha_{g}).$
Note that $A_{g,r,d}$ is independent of $d$, so we denote it $A_{g,r}$ henceforth. When applied to $k=\mathbb{C}$, Mellit’s result proved a conjectural formula of Hausel and Rodriguez-Villegas for the $\mathbb{Q}$-Betti numbers of $M_{r}^{d}$. Theorem 1.1 can be thought of as a Hausel Rodriguez-Villegas style formula for the $\mathbb{Z}_{2}$-Betti number of the moduli space of real Higgs bundles over a real curve.
Our proof of Theorem 1.1 can be outlined as follows. Fedorov-Soibelman-Soibelman [FSS] proved a motivic version of Schiffmann’s formula which, when $k$ is a field of characteristic zero, identifies the class $M_{r}^{d}]$ in $\overline{Mot}(k)$, the dimensional completion of the Grothendieck ring of Artin stacks over $k$. Mellit ([M] §6), identifies their class $M_{r}^{d}]$ as the image of $q^{(g-1)r^{2}+1}A_{r,d}$ under a homomorphism
(1.2) $ev_{\Sigma}:R_{g}\rightarrow\overline{Mot}(k),$
where $R_{g}$ is the ring of Laurent polynomials in $q,\alpha_{1}^{\pm 1},...,\alpha_{g}^{\pm 1}$ which are invariant under permutations of $\alpha_{i}$ and under $\alpha_{i}\mapsto q\alpha_{i}^{-1}$. This $ev_{\Sigma}$ sends $q$ to the Tate class $\mathbb{L}$ and the extension to $R_{g}[[z]]\rightarrow\overline{Mot}(k)[[z]]$ sends the formal zeta function of the curve to the motivic zeta function of $\Sigma$ :
(1.3) $ev_{\Sigma}\left(\frac{\prod_{i=1}^{g}(1-\alpha_{i}z)(1-\alpha_{i}^{-1}qz)}{(1-z)(1-qz)}\right)=\sum_{n=0}^{\infty}z^{n}\Sigma^{(n)}=Z_{\Sigma}(z).$
When $k=\mathbb{R}$, one can define a virtual Poincaré polynomial homomorphism
$P_{t}^{vir}:\overline{Mot}(\mathbb{R})\rightarrow\mathbb{Z}((t^{-1})),$
which sends the class $[X]$ of a smooth, projective variety $X$ to the Poincaré polynomial $P_{t}^{vir}([X])=P_{t}(X(\mathbb{R}))$.
To calculate $P_{t}^{vir}([M_{r}^{d}])$, it only remains to determine the composition $P_{t}^{vir}\circ ev_{\Sigma}$, which boils down to calculating $P_{t}^{vir}(Z_{\Sigma}(z))$. We carry this out in §3. Then in §4 we prove that the virtual Poincaré polynomial $M_{r}^{d}$ agrees with its topological Poincaré polynomial, completing the proof of Theorem 1.1. In section 5 we show that the Poincaré polynomial $P_{t}(M_{r}^{d})$ is divisible by $P_{t}(Pic^{0}(\Sigma))$, which has the surprising implication that average value of the Poincaré polynomial of the connected components of $M_{r}^{d}$ is a polynomial with integer coefficients.
Acknowledgements
Thank you to Oscar Garcia-Prada and to Ajneet Dhillon for suggesting that I use motivic formulas to calculate Betti numbers. Thanks also to Oscar Garcia-Prada and Jochen Heinloth who shared Maple code to calculate $M^{d}_{4}]$ in terms of the motivic formula derived in [GHS]. Indeed the original plan for the current paper was to work from [GHS], and though their formula was later superseded by [FSS], it remained very helpful for confirming calculations.
2. The Grothendieck Ring of Stacks
We denote by $K_{0}(Var_{k})$ the Grothendieck ring of varieties over the field $k$. This is the free abelian group generated by isomorphism classes $[X]$ of quasi-projective varieties over $X/k$ modulo the relation $X]=[X\setminus Y]+[Y]$ where $Y\subseteq X$ is a closed subvariety. Multiplication is determined by the rule $X]\cdot[Y]=[X\times_{k}Y]$. Denote by $\mathbb{L}:=[\mathbb{A}^{1}]\in K_{0}(Var_{k})$, called the Lefschetz element.
Similarly, denote by $Mot(k)$ the Grothendieck ring of Artin stacks over $k$. This is the free abelian group generated by isomorphism classes $\mathcal{X}]$ of algebraic stacks $\mathcal{X}/k$ of finite type with affine stabilizers, modulo the relation $\mathcal{X}]=[\mathcal{X}\setminus\mathcal{Y}]+[\mathcal{Y}]$ if $\mathcal{Y}$ is a closed substack of $\mathcal{X}$ and the relation $\mathcal{E}]=[\mathcal{X}][\mathbb{A}^{n}]$ if $\mathcal{E}\rightarrow\mathcal{X}$ is a vector bundle of rank $n$ (this second relation is automatic for $K_{0}(Var_{k})$). Multiplication is determined by $\mathcal{X}][\mathcal{Y}]=[\mathcal{X}\times\mathcal{Y}]$.
There is a natural homomorphism $K_{0}(Var_{k})\rightarrow Mot(k)$ sending $[X]$ to $[X]$. This descends to an isomorphism $K_{0}^{\prime}(Var_{k})\cong Mot(k)$ where $K_{0}^{\prime}(Var_{k})$ is the localization of $K_{0}(Var_{k})$ obtained by inverting $\mathbb{L}$ and $(\mathbb{L}^{n}-1)$ for all $n=1,2,3,$... ([E] Thm 1.2).
We have a filtration $F^{1}\subset F^{2}\subset...\subset Mot(k)$ where $F^{n}$ is the subgroup generated by classes $\mathcal{X}]$ with $\dim(\mathcal{X})\leq-n$. The dimensional completion $\overline{Mot}(k)$ is the completion of $Mot(k)$ under this filtration.
For a real projective variety $X/\mathbb{R}$, denote by
$P_{t}(X):=\sum_{i=0}^{\infty}(-1)^{i}b_{i}(X)t^{i}$
where $b_{i}(X)=\dim H^{i}_{cpt}(X(\mathbb{R});\mathbb{Z}_{2})$ is the dimension of the $i$th degree compactly supported cohomology with $\mathbb{Z}_{2}$ coefficients of $X(\mathbb{R})$ endowed with analytic topology.
Proposition 2.1.
There exists a unique motivic measure $P_{t}^{vir}:K_{0}(Var_{\mathbb{R}})\rightarrow\mathbb{Z}[t]$ such that for all smooth projective $X/\mathbb{R}$ we have $P_{t}^{vir}([X])=P_{t}(X)$. This extends naturally to a ring homomorphism
$P_{t}^{vir}:\overline{Mot}(\mathbb{R})\rightarrow\mathbb{Z}[t][[t^{-1}]]=\mathbb{Z}((t^{-1})).$
Proof.
The homomorphism
$P_{t}^{vir}:K_{0}(Var_{\mathbb{R}})\rightarrow\mathbb{Z}[t]\subset\mathbb{Z}((t^{-1}))$
is constructed by McCrory and Parusinski [MP]. Since $P_{t}^{vir}(\mathbb{L})=-t$ and $P_{t}^{vir}(\mathbb{L}^{n}-1))=(-t)^{n}-1$ are both invertible in $\mathbb{Z}((t^{-1}))$, with $((-t)^{n}-1)^{-1}=(-t)^{-n}(\sum_{k=0}^{\infty}(-t)^{-kn})$, this extends to a homomorphism
$P_{t}^{vir}:Mot(\mathbb{R})\rightarrow\mathbb{Z}((t^{-1})),$
which extends naturally to the dimensional completion. ∎
The zeta function of a variety $X$ is given by the formal power series
$Z_{X}(z):=\sum_{n=0}^{\infty}[X^{(n)}]z^{n}\in Mot(k)[[z]],$
where $X^{(n)}=X^{n}/S_{n}$ denotes the $n$-fold symmetric product of $X$. The zeta function is multiplicative in the sense that for closed $Y\subseteq X$,
$Z(X,z)=Z(Y,z)Z(X\setminus Y,z).$
If $\Sigma/k$ is a smooth projective curve of genus $g$ and $\Sigma(k)\neq\emptyset$, then a theorem of Kapranov ([K] Theorem 1.1.9) says that
$Z_{\Sigma}(z)=\frac{P(z)}{(1-z)(1-\mathbb{L}z)}$
where $P(z)$ is a polynomial of degree $2g$ such that $Z_{\Sigma}(1/\mathbb{L}z)=\mathbb{L}^{1-g}x^{2-2g}Z_{\Sigma}(z)$. The homomorphism (1.2) is determined by the property that $\prod_{i=1}^{g}(1-\alpha_{i}z)(1-\alpha_{i}^{-1}qz)$ is sent to $P(z)$ in some algebraic extension of $\overline{Mot}(k)$
3. Betti numbers of $\Sigma^{(n)}(\mathbb{R})$
Suppose that $\Sigma$ is a smooth, geometrically connected, projective curve of genus $g$ over $\mathbb{R}$. For $n\geq 0$ denote by $\Sigma^{(n)}$ its $n$-th symmetric product. The set of complex points $\Sigma^{(n)}(\mathbb{C})$ can be identified with the orbit space
$Sym^{n}(\Sigma(\mathbb{C})):=\Sigma(\mathbb{C})^{n}/S_{n},$
while $\Sigma^{(n)}(\mathbb{R})\subset\Sigma^{(n)}(\mathbb{C})$ can be identified with $Gal(\mathbb{C}/\mathbb{R})$ fixed points.
If $\Sigma(\mathbb{R})$ is empty, then
$\Sigma^{(n)}(\mathbb{R})\cong\begin{cases}\emptyset&\text{if $n$ is odd}\\ Sym^{n/2}(N_{g})&\text{if $n$ is even}\end{cases}$
where $N_{g}$ is a non-orientable surface double covered by $\Sigma(\mathbb{C})$. We abusively write $P_{t}(X):=\sum_{k}(-t)^{k}H_{cpt}^{k}(X;\mathbb{Z}_{2})$. By Macdonald’s formula [Mac],
(3.1) $\sum_{m=0}^{\infty}P_{t}(Sym^{m}(N_{g}))z^{m}=\frac{(1-tz)^{g+1}}{(1-z)(1-t^{2}z)},$
so
$P_{t}^{vir}(Z_{\Sigma}(z)):=\sum_{n=0}^{\infty}P_{t}(\Sigma^{(n)})z^{n}=\frac{(1-tz^{2})^{g+1}}{(1-z^{2})(1-t^{2}z^{2})}.$
When $\Sigma(\mathbb{R})\neq\emptyset$, $P_{t}(\Sigma^{(n)})$ was calculated in [B].
Proposition 3.1.
If $\Sigma(\mathbb{R})$ is non-empty, with $b+1$ many connected components, then
(3.2) $P_{t}^{vir}(Z_{\Sigma}(z))=\sum_{n=0}^{\infty}P_{t}(\Sigma^{(n)})z^{n}=\frac{(1-tz^{2})^{g-b}(1+z)^{b}(1-tz)^{b}}{(1-z)(1+tz)}.$
Proof.
Suppose first that $\Sigma(\mathbb{C})\setminus\Sigma(\mathbb{R})$ is connected. For each pair of integers $k,s\geq 0$ such that $k+2s=n$, $\Sigma^{(n)}(\mathbb{R})$ has $b+1\choose k$} connected components with Poincaré polynomial isomorphic to $(1-t)^{k}P_{t}(Sym^{s}(N_{g-k}))$. Therefore, setting $a=b+1$, we have
$\sum_{n=0}^{\infty}P_{t}(\Sigma^{(n)})z^{m}$ $=$ $\sum_{0\leq k\leq a}\sum_{s\geq 0}z^{k+2s}{a\choose k}(1-t)^{k}P_{t}(Sym^{s}(N_{g-k}))$
[…]
$=$ $\left((z-zt)+(1-tz^{2})\right)^{a}\frac{(1-tz^{2})^{g-a+1}}{(1-z^{2})(1-t^{2}z^{2})}$
$=$ $\frac{(1-tz^{2})^{g-a+1}(1+z)^{a-1}(1-tz)^{a-1}}{(1-z)(1+tz)}$
In case $\Sigma(\mathbb{C})\setminus\Sigma(\mathbb{R})$ is not connected, the calculation is the same except that whenever $k=a$, the corresponding component has Poincaré polynomial $(1-t)^{b+1}P_{t}(Sym^{s}(\Sigma_{(g-b)/2}))$, where $\Sigma_{g}$ is an orientable surface of genus $g$. But this equals $(1-t)^{b+1}P_{t}(Sym^{s}(N_{g-b-1}))$ ([B] Remark 3.14) so the same formula holds. ∎
Corollary 3.2.
With hypotheses of Theorem 1.1 we have
$P_{t}^{vir}([M_{r}^{d}])=(-t)^{(g-1)r^{2}+1}A_{g,r}(-t,-t^{\frac{1}{2}},...,-t^{\frac{1}{2}},-1,...,-1).$
Proof.
Substituting q=tq=-t, $\alpha_{1}=...=\alpha_{g-b}=-t^{\frac{1}{2}}$ and $\alpha_{g-b+1}=...=\alpha_{g}=-1$ into the formal zeta function
$\frac{\prod_{i=1}^{g}(1-\alpha_{i}z)(1-\alpha_{i}^{-1}qz)}{(1-z)(1-qz)}$
yields (3.2). Apply (1.3). ∎
4. On purity of real semi-projective varieties
A semi-projective variety $Y$ over $\mathbb{C}$ is a smooth, quasiprojective variety equipped with $\mathbb{C}^{\times}$-action such that the fixed point set $Y^{\mathbb{C}^{\times}}$ is proper and for ever $y\in Y$ the limit $lim_{\lambda\rightarrow 0}\lambda y$ exists as $\lambda\in\mathbb{C}^{\times}$ tends to zero. Mixed Hodge structures of such varieties were studied in Hausel and Rodriguez-Villegas [HRV13] .
A semi-projective variety $X$ over $\mathbb{R}$ is a real variety equipped with an action of $\mathbb{R}^{\times}$ such that the base change $X\times_{R}\mathbb{C}$ is semi-projective over $\mathbb{C}$. When $\Sigma$ is defined over $\mathbb{R}$, $M_{r}^{d}$ is a semi-projective variety, with respect the $\mathbb{C}^{\times}$ action of scaling the Higgs field.
Proposition 4.1.
If $X$ is semi-projective over $\mathbb{R}$ then
$P_{t}^{vir}(X)=P_{t}(X).$
Proof.
Under these conditions $X^{\mathbb{C}^{\times}}=\coprod_{i\in I}F_{i}$ where the $F_{i}$ are non-singular, projective over $\mathbb{R}$, and we have a Bialycki-Birula stratification $X=\bigcup_{i\in I}U_{i}$ where
$U_{i}(\mathbb{C}):=\{x\in X(\mathbb{C})|lim_{\lambda\mapsto 0}\lambda x\in F_{i}\},$
and each $U_{i}$ is a locally closed subvarity of $X$, isomorphic to a vector bundle of some rank $d_{i}$ over $F_{i}$. It follows that
$P_{t}^{vir}(X)=\sum_{i\in I}P_{t}^{vir}(U_{i})=\sum_{i\in I}P_{t}^{vir}(F_{i})P_{t}^{vir}(\mathbb{A}^{d_{i}})=\sum_{i\in I}P_{t}(F_{i})(-t)^{d_{i}}.$
On the other hand, by a result essentially due to Duistermaat [D] (see also [BGH]), we know that $P_{t}(X)=\sum_{i}P_{t}(U_{i})$, and $U_{i}(\mathbb{R})$ deformation retracts onto $F_{i}(\mathbb{R})$ and has relative dimension $d_{i}$, so by Poincaré duality, $P_{t}(U_{i})=P_{t}(F_{i})(-t)^{d_{i}}$, completing the proof. ∎
5. Average Betti numbers
Our goal in this section is to prove the following somewhat surprising result.
Proposition 5.1.
Suppose that $\Sigma$ is a smooth projective, geometrically connected curve of genus $g$ defined over $\mathbb{R}$, and $\Sigma(\mathbb{R})$ has $b+1>0$ connected components. Then the Poincaré polynomial $P_{t}(M_{r}^{d})$ is divisible by $2^{b}(1-t)^{g}$.
Proof.
Recall (1.1):
(5.1) $\sum_{r=1}^{\infty}H_{g,r}T^{r}=(z-1)(1-q)\operatorname{Log}\sum_{\mu\in\mathcal{P}}T^{|\mu|}\mathcal{H}_{\mu}.$
Every non-empty Young diagram $\mu$ includes at least one box with both arm and leg length zero, so $\mathcal{H}_{\mu}$ contains a factor of $\prod_{i=1}^{g}(z-\alpha_{i})(1-\alpha_{i}^{-1}q)$ in its numerator. Using the combinatorics of the plethystic logarithm (see e.g. [HRV08] Theorem 3.5.2.), it follows that every $H_{g,r}$ is divisible by $\prod_{i=1}^{g}(z-\alpha_{i})(1-\alpha_{i}^{-1}q)$ in the localized ring $S^{-1}\mathbb{Q}[q,z,\alpha_{1}^{\pm 1},...,\alpha_{g}^{\pm 1}]$ where $S$ is the multiplicative set generated by $\{(z^{a+1}-q^{l}),(z^{a}-q^{l+1})|a,b\geq 0\}$. We know
$A_{g,r}(q,\alpha_{1},...,\alpha_{g})=H_{g,r}(q,1,\alpha_{1},...,\alpha_{g})\in\mathbb{Z}[q,\alpha_{1}^{\pm 1},...,\alpha_{g}^{\pm 1}]$
[…]
$2^{b}(1-t)^{g}=(1+t^{\frac{1}{2}})^{g-b}(1-t^{\frac{1}{2}})^{g-b}(1+1)^{b}(1-t)^{b}$
in $\mathbb{Z}[t^{\pm\frac{1}{2}}]$, hence also in $\mathbb{Z}[t]$. ∎
When the rank $r$ is odd, Proposition 5.1 can be proven using more elementary reasoning. Choose a real line bundle $L$ over $\Sigma$ of degree $d$, and let $M_{r}^{L}$ be the moduli space of stable Higgs bundles of rank $r$ with determinant $L$. The morphism $M_{r}^{L}\times Pic(\Sigma)^{0}\rightarrow M_{r}^{d}$ sending the Higgs bundle $(E,\varphi)$ to $(E\otimes L,\varphi\otimes Id_{L})$ descends to an isomorphism $M_{r}^{L}\times_{\Gamma_{r}}Pic^{0}(\Sigma)$ where $\Gamma_{r}\leq Pic^{0}(\Sigma)$ is the $r$-torsion subgroup scheme.
When $r$ is odd, this determines a homeomorphism
(5.2) $M_{r}^{L}(\mathbb{R})\times_{\Gamma_{r}(\mathbb{R})}Pic^{0}(\Sigma)(\mathbb{R})\cong M_{r}^{d}(\mathbb{R}).$
Here $M_{r}^{L}(\mathbb{R})$ is connected, $Pic^{0}(\Sigma)(\mathbb{R})\cong\mathbb{Z}_{2}^{b}\times(S^{1})^{g}$ as a Lie group, and $\Gamma_{r}$ lies in the identity component of $Pic^{0}(\Sigma)(\mathbb{R})$, which implies that $M_{r}^{d}(\mathbb{R})$ has $2^{b}$ pairwise isomorphic connected components. Since $|\Gamma_{r}|$ and $|\mathbb{Z}_{2}|$ are relatively prime,
$H^{*}(M_{r}^{d}(\mathbb{R}),\mathbb{Z}_{2})$ $\cong$ $\left(H^{*}(M_{r}^{L}(\mathbb{R});\mathbb{Z}_{2})\otimes H^{*}(Pic^{0}(\Sigma)(\mathbb{R});\mathbb{Z}_{2})\right)^{\Gamma_{r}(\mathbb{R})}$
$\cong$ $H^{*}(M_{r}^{L}(\mathbb{R});\mathbb{Z}_{2})^{\Gamma_{r}(\mathbb{R})}\otimes H^{*}(Pic^{0}(\Sigma)(\mathbb{R});\mathbb{Z}_{2}),$
so $P_{t}(M_{r}^{d})$ must be divisible by $P_{t}(Pic^{0}(\Sigma))=2^{b}(1-t)^{g}$.
This argument fails when $r$ is even. In this case $\Gamma_{r}(\mathbb{R})$ intersects every connected component of $Pic^{0}(\mathbb{R})$, so (5.2) maps to only one of the $2^{b}$ components of $M_{r}^{d}(\mathbb{R})$. Indeed, the connected components of $M_{r}^{d}(\mathbb{R})$ have different Poincaré polynomials in general. For example (see [B] §5.2), when $g=b=2$, $M_{2}^{1}(\mathbb{R})$ has four components, three of which have Poincaré polynomial $t^{5}(1-t)^{2}(5t^{3}-7t^{2}+3t-1)$ and one of which has Poincaré polynomial $t^{5}(1-t)^{2}(t^{3}-3t^{2}+3t-1)$. Adding these together gives $4t^{5}(1-t)^{2}(4t^{3}-6t^{2}+3t-1)$. A priori, it is not at all obvious that the average value of the Poincaré polynomials of the connected components should have integer coefficients.